Section 15

Numerical Analysis and Scientific Computing

In case of several authors, Invited Speakers are marked with a ^{∗}.

Gregory Beylkin: On Multiresolution Methods in Numerical Analysis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 481

P. Deift^{∗}, T. Kriecherbauer, K. T-R McLaughlin, S.

Venakides and X. Zhou: Uniform Asymptotics for Orthogonal

Polynomials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 491

Bjorn Engquist: Wavelet Based Numerical Homogenization . . . . . . . . III 503

Hisashi Okamoto: A Study of Bifurcation of Kolmogorov Flows with an Emphasis on the Singular Limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 513

Jan-Olov Strömberg: Computation with Wavelets in Higher Dimensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 523

Lloyd N. Trefethen^{∗} and Tobin A. Driscoll:

Schwarz–Christoffel Mapping in the Computer Era . . . . . . . . . . . . . . . . . III 533

Documenta Mathematica · Extra Volume ICM 1998 · III ·

480

Documenta Mathematica · Extra Volume ICM 1998 · III ·